OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x - 1)?

OpenStudy (anonymous):

5(x+2) >= 7x + 2(x-1) 5x+10 >= 7x + 2x -2 Then solve: 5x >= 9x -2+10 5x -9x >= 8 -4x >= 8 (-4x>=8)/-4 x >= -2 The answer is x is greater than or equal to -2.